OpenStudy (anonymous):

\[\frac{1}{2}\int_{b-1/2}^{b+1/2}\frac{1}{\sqrt{|x|}}dx=\frac{1}{2}\int_{b-1/2}^{0}\frac{1}{\sqrt{|x|}}dx+\frac{1}{2}\int_{0}^{b+1/2}\frac{1}{\sqrt{|x|}}dx=\sqrt{1/2-b}+\sqrt{1/2+b}\]

OpenStudy (anonymous):

How can I format this to fit the window?

OpenStudy (anonymous):

\[\frac{1}{2}\int_{b-1/2}^{b+1/2}\frac{1}{\sqrt{|x|}}dx=\] \[\frac{1}{2}\int_{b-1/2}^{0}\frac{1}{\sqrt{|x|}}dx+\frac{1}{2}\int_{0}^{b+1/2}\frac{1}{\sqrt{|x|}}dx=\] \[\sqrt{1/2-b}+\sqrt{1/2+b}\]

OpenStudy (theviper):

Yes I too have the same question :/ @satellite73 @sasogeek @ParthKohli Plz help :)

OpenStudy (anonymous):

I simply broke the equation up in three parts manually :-)

OpenStudy (kinggeorge):

\[\begin{aligned} \frac{1}{2}\int_{b-1/2}^{b+1/2}\frac{1}{\sqrt{|x|}}dx &= \frac{1}{2}\int_{b-1/2}^{0}\frac{1}{\sqrt{|x|}}dx+\frac{1}{2}\int_{0}^{b+1/2}\frac{1}{\sqrt{|x|}}dx \\ &=\sqrt{1/2-b}+\sqrt{1/2+b} \end{aligned}\]There's no way to get it to format automatically. However, you can use the environment `\begin{aligned}` and `\end{aligned}` to format lines in regards to one another and make it look pretty nice. The code I used to make the above math was `\begin{aligned}` `\frac{1}{2}\int_{b-1/2}^{b+1/2}\frac{1}{\sqrt{|x|}}dx` `&= \frac{1}{2}\int_{b-1/2}^{0}\frac{1}{\sqrt{|x|}}dx` `+\frac{1}{2}\int_{0}^{b+1/2}\frac{1}{\sqrt{|x|}}dx \\` `&=\sqrt{1/2-b}+\sqrt{1/2+b}` `\end{aligned}` "\\" marks a new line, and the "&" signs are automatically lined up.

OpenStudy (unklerhaukus):

\[\begin{aligned} \frac 12\int\limits_{b-1/2}^{b+1/2}\frac{dx}{\sqrt{|x|}} &= \frac12\int\limits_{b-1/2}^{0}\frac{dx}{\sqrt{|x|}} +\frac12\int\limits_{0}^{b+1/2}\frac{dx}{\sqrt{|x|}}\\ &=\sqrt{1/2-b}+\sqrt{1/2+b} \end{aligned}\] \[\begin{aligned} \frac 12\int\limits_{b-1/2}^{b+1/2}\frac{dx}{\sqrt{|x|}} &= \frac12\int\limits_{b-1/2}^{0}\frac{dx}{\sqrt{|x|}} +\frac12\int\limits_{0}^{b+1/2}\frac{dx}{\sqrt{|x|}}\\ &=\sqrt{1/2-b}+\sqrt{1/2+b} \end{aligned }\]

Parth (parthkohli):

@KingGeorge ``` ``` create multi-line here ``` ```

OpenStudy (unklerhaukus):

``` \begin{aligned} \frac 12\int\limits_{b-1/2}^{b+1/2}\frac{dx}{\sqrt{|x|}} &=\frac 12\int\limits_{b-1/2}^{0}\frac{dx{\sqrt{|x|}} +\frac12\int\limits_{0}^{b+1/2}\frac{dx}{\sqrt{|x|}}\\ &=\sqrt{1/2-b}+\sqrt{1/2+b}\end{aligned} ``` ah, pretty colours

OpenStudy (anonymous):

\(\def\HappyBirthdayAbhisar{\small\sf\color{teal}{\mbox{Happy birthday, to the best friend a guy could ever ask for nobody could ever replace you}\\\mbox{you truly are amzing, kind, helpful, cheerful, funny, fantastic, good listener, problem solver,}\\\mbox{fight resolver, great moderator, kind, caring, respectful. and so much more that i cant even list!}\\\mbox{Never change, always be yourself... were always here for you and all of us greatly appreciate}\\\mbox{what you do on a daily basis for all of us even if your extremly annoyed you never let it bother you}\\\mbox{you power through it and help us in our time of need. You truly are the BEST friend a gut could }\\\mbox{EVER ask for, never stop being you abhisar because you are great... Happy Birthday My Friend!}}\\\scriptsize\bbox[1pt,#ff002f]{\bbox[1pt,#ff3f00]{\bbox[0.1pt,#ff6c00]{\bbox [0.1pt,#ff8200]{\bbox[0.1pt,#ff8700]{\bbox[0.1pt,#ff9900]{\bbox[0.1pt,#ff9f00]{\bbox [0.1pt,#ffad00]{\bbox[0.1pt,#ffb700]{\bbox[0.1pt,#ffc100]{\bbox[0.1pt,#ffcf00]{\bbox [0.1pt,#ffd900]{\bbox[0.1pt,#ffe200]{\bbox[0.1pt,#ffe500]{\bbox[0.1pt,#ffec00]{\bbox [0.1pt,#ffef00]{\bbox[0.1pt,#fff400]{\bbox[0.1pt,#fff800]{\bbox[0.1pt,#fffb00]{\bbox [0.1pt,#f9ff00]{\bbox[0.1pt,#f0ff00]{\bbox[0.1pt,#e7ff00]{\bbox[0.1pt,#e0ff00]{\bbox [0.1pt,#d8ff00]{\bbox[0.1pt,#d1ff00]{\bbox[0.1pt,#c8ff00]{\bbox[0.1pt,#bfff00]{\bbox [0.1pt,#b9ff00]{\bbox[0.1pt,#b0ff00]{\bbox[0.1pt,#a8ff00]{\bbox[0.1pt,#9cff00]{\bbox [0.1pt,#90ff00]{\bbox[0.1pt,#85ff00]{\bbox[0.1pt,#77ff00]{\bbox[0.1pt,#66ff00]{\bbox [0.1pt,#51ff00]{\bbox[0.1pt,#42ff00]{\bbox[0.1pt,#34ff00]{\bbox[0.1pt,#1fff00]{\bbox [0.1pt,#0cff00]{\bbox[0.1pt,#02ff00]{\bbox[0.1pt,#00ff11]{\bbox[0.1pt,#00ff28]{\bbox [0.1pt,#00ff36]{\bbox[0.1pt,#00ff46]{\bbox[0.1pt,#00ff58]{\bbox[0.1pt,#00ff6d]{\bbox [0.1pt,#00ff83]{\bbox[0.1pt,#00ff96]{\bbox[0.1pt,#00ffa4]{\bbox[0.1pt,#00ffa8]{\bbox [0.1pt,#00ffae]{\bbox[0.1pt,#00ffb9]{\bbox[0.1pt,#00ffc5]{\bbox[0.1pt,#00ffd2]{\bbox [0.1pt,#00ffe0]{\bbox[0.1pt,#00ffeb]{\bbox[0.1pt,#00fffa]{\bbox[0.1pt,#00f7ff]{\bbox [0.1pt,#00ebff]{\bbox[0.1pt,#00e0ff]{\bbox[0.1pt,#00d3ff]{\bbox[0.1pt,#00c8ff]{\bbox [0.1pt,#00b2ff]{\bbox[0.1pt,#009fff]{\bbox[0.1pt,#008aff]{\bbox[0.1pt,#0073ff]{\bbox [0.1pt,#005dff]{\bbox[0.1pt,#0044ff]{\bbox[0.1pt,#002bff]{\bbox[0.1pt,#001bff]{\bbox [0.1pt,#0000ff]{\bbox[0.1pt,#1600ff]{\bbox[0.1pt,#2e00ff]{\bbox[0.1pt,#4400ff]{\bbox [0.1pt,#5b00ff]{\bbox[0.1pt,#7600ff]{\bbox[0.1pt,#8800ff]{\bbox[0.1pt,#9d00ff]{\bbox [0.1pt,#b500ff]{\bbox[0.1pt,#c700ff]{\bbox[0.1pt,#d600ff]{\bbox[0.1pt,#ee00ff]{\bbox [0.1pt,#fb00ff]{\bbox[0.1pt,#ff00e5]{\bbox[0.1pt,#ff00c6]{\bbox[0.1pt,#ff00a7]{\bbox [0.1pt,#ff0087]{\bbox[0.1pt,#ff0061]{\bbox[0.1pt,#ff0045]{\bbox[0.1pt,#ff0031]{{\Huge \rlap{\bbox[2.5pt, orange ,border:3px solid red ]{\small\color{darkred}{\bf \mbox{Happiest of Birthdays to a truly great friend have a great day Abhisar!}}}}\bbox[2.5pt, orange ,border:0px solid red ]{\small\color {red}{\bf{\hspace{3.726pt}Happiest\hspace{2.1pt}of\hspace{1.5pt}Birthdays\hspace{2pt}to\hspace{2pt}a\hspace{2pt}truly\hspace{2pt}great\hspace{2pt} friend\hspace{2pt}have\hspace{2pt}a\hspace{2pt}great\hspace{2pt}day\hspace{2pt} Abhisar!}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}} }}}}}}}}\HappyBirthdayAbhisar\)

jaynator495 (jaynator495):

\(\newcommand{HappyBirthdayAbhisar}{\small\sf\color{teal}{\mbox{Happy birthday, to the best friend a guy could ever ask for nobody could ever replace you}\\\mbox{you truly are amzing, kind, helpful, cheerful, funny, fantastic, good listener, problem solver,}\\\mbox{fight resolver, great moderator, kind, caring, respectful. and so much more that i cant even list!}\\\mbox{Never change, always be yourself... were always here for you and all of us greatly appreciate}\\\mbox{what you do on a daily basis for all of us even if your extremly annoyed you never let it bother you}\\\mbox{you power through it and help us in our time of need. You truly are the BEST friend a gut could }\\\mbox{EVER ask for, never stop being you abhisar because you are great... Happy Birthday My Friend!}}\\\scriptsize\bbox[1pt,#ff002f]{\bbox[1pt,#ff3f00]{\bbox[0.1pt,#ff6c00]{\bbox[0.1pt,#ff8200]{\bbox [0.1pt,#ff8700]{\bbox[0.1pt,#ff9900]{\bbox[0.1pt,#ff9f00]{\bbox[0.1pt,#ffad00]{\bbox[0.1pt,#ffb700]{\bbox[0.1pt,#ffc100]{\bbox[0.1pt,#ffcf00]{\bbox[0.1pt,#ffd900]{\bbox[0.1pt,#ffe200]{\bbox[0.1pt,#ffe500]{\bbox [0.1pt,#ffec00]{\bbox[0.1pt,#ffef00]{\bbox[0.1pt,#fff400]{\bbox[0.1pt,#fff800]{\bbox[0.1pt,#fffb00]{\bbox[0.1pt,#f9ff00]{\bbox[0.1pt,#f0ff00]{\bbox[0.1pt,#e7ff00]{\bbox[0.1pt,#e0ff00]{\bbox[0.1pt,#d8ff00]{\bbox [0.1pt,#d1ff00]{\bbox[0.1pt,#c8ff00]{\bbox[0.1pt,#bfff00]{\bbox[0.1pt,#b9ff00]{\bbox[0.1pt,#b0ff00]{\bbox[0.1pt,#a8ff00]{\bbox[0.1pt,#9cff00]{\bbox[0.1pt,#90ff00]{\bbox[0.1pt,#85ff00]{\bbox[0.1pt,#77ff00]{\bbox [0.1pt,#66ff00]{\bbox[0.1pt,#51ff00]{\bbox[0.1pt,#42ff00]{\bbox[0.1pt,#34ff00]{\bbox[0.1pt,#1fff00]{\bbox[0.1pt,#0cff00]{\bbox[0.1pt,#02ff00]{\bbox[0.1pt,#00ff11]{\bbox[0.1pt,#00ff28]{\bbox[0.1pt,#00ff36]{\bbox [0.1pt,#00ff46]{\bbox[0.1pt,#00ff58]{\bbox[0.1pt,#00ff6d]{\bbox[0.1pt,#00ff83]{\bbox[0.1pt,#00ff96]{\bbox[0.1pt,#00ffa4]{\bbox[0.1pt,#00ffa8]{\bbox[0.1pt,#00ffae]{\bbox[0.1pt,#00ffb9]{\bbox[0.1pt,#00ffc5]{\bbox [0.1pt,#00ffd2]{\bbox[0.1pt,#00ffe0]{\bbox[0.1pt,#00ffeb]{\bbox[0.1pt,#00fffa]{\bbox[0.1pt,#00f7ff]{\bbox[0.1pt,#00ebff]{\bbox[0.1pt,#00e0ff]{\bbox[0.1pt,#00d3ff]{\bbox[0.1pt,#00c8ff]{\bbox[0.1pt,#00b2ff]{\bbox [0.1pt,#009fff]{\bbox[0.1pt,#008aff]{\bbox[0.1pt,#0073ff]{\bbox[0.1pt,#005dff]{\bbox[0.1pt,#0044ff]{\bbox[0.1pt,#002bff]{\bbox[0.1pt,#001bff]{\bbox[0.1pt,#0000ff]{\bbox[0.1pt,#1600ff]{\bbox[0.1pt,#2e00ff]{\bbox [0.1pt,#4400ff]{\bbox[0.1pt,#5b00ff]{\bbox[0.1pt,#7600ff]{\bbox[0.1pt,#8800ff]{\bbox[0.1pt,#9d00ff]{\bbox[0.1pt,#b500ff]{\bbox[0.1pt,#c700ff]{\bbox[0.1pt,#d600ff]{\bbox[0.1pt,#ee00ff]{\bbox[0.1pt,#fb00ff]{\bbox [0.1pt,#ff00e5]{\bbox[0.1pt,#ff00c6]{\bbox[0.1pt,#ff00a7]{\bbox[0.1pt,#ff0087]{\bbox[0.1pt,#ff0061]{\bbox[0.1pt,#ff0045]{\bbox[0.1pt,#ff0031]{{\Huge\rlap{\bbox[2.5pt, orange ,border:3px solid red ]{\small\color {darkred}{\bf\mbox{Happiest of Birthdays to a truly great friend have a great day Abhisar!}}}}\bbox[2.5pt, orange ,border:0px solid red ]{\small\color{red}{\bf{\hspace{3.726pt}Happiest\hspace{2.1pt}of\hspace{1.5pt} Birthdays\hspace{2pt}to\hspace{2pt}a\hspace{2pt}truly\hspace{2pt}great\hspace{2pt}friend\hspace{2pt}have\hspace{2pt}a\hspace{2pt}great\hspace{2pt}day\hspace{2pt} Abhisar!}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}}{ \begin{matrix}{\HappyBirthdayAbhisar} \end{matrix}}\)